MrBibbers:

Which expression is a difference of squares with a factor of 5x − 8? 25x2 − 16 25x2 + 16 25x2 − 64 25x2 + 64

Ultrilliam:

a2 - b2 = (a-b)(a+b) (5x-8)(5x+8) = 25x2 - 64 25x2 - 64